OpenStudy (anonymous):

Given the equation 7x + 2 = 2 + 5x – 6, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 7x + 2 = 2 – 6 + 5x. Then, use the ____________________ to group the like terms. This gives the equation 7x + 2 = (2 – 6) + 5x.

OpenStudy (anonymous):

Commutative Property of Addition Associative Property of Addition Distributive Property Addition Property of Equality

OpenStudy (anonymous):

i think it is equality one but idk

OpenStudy (anonymous):

what a weird question i think they are looking for "associative"

OpenStudy (radar):

7x-5x=-6 x(7-5)=-6 2x=-6 x=-3 I'm not sure of the objective of this exercise.

OpenStudy (anonymous):

compare \[7x + 2 = 2 – 6 + 5x\] with \[7x + 2 = (2 – 6) + 5x\] and we see that the only difference is the introduction of parentheses associative law says \((a+b)+c=a+(b+c)\) but the first expression contains no parentheses, so it is a bit confusing

OpenStudy (anonymous):

ya i did the math but was not sure what 'property' they were

OpenStudy (anonymous):

associative law says parentheses are not necessary, so i think that is what the author of this rather idiosyncratic question is looking for